OpenStudy (anonymous):

Write as a single logarithm 3log(x-2) - 2log(x-1)

OpenStudy (anonymous):

Do you know any of the log rules? If not I'll introduce them =)

OpenStudy (anonymous):

I know some of them, I just dont fully understand them

OpenStudy (anonymous):

\[a \log x=\log x^a\]

OpenStudy (anonymous):

\[\log m-\log n=\log \frac{ m }{ n }\]

OpenStudy (anonymous):

Yup, @surjithayer got here before me. Can you see how to use that in your question? Once you have that bit sorted we can look at merging them into one logarithm.

OpenStudy (anonymous):

so is it log(x-2)^3 / log(x-1)^2?

OpenStudy (anonymous):

or log (x-2)^3 over (x-1)^2

OpenStudy (jdoe0001):

http://www.chilimath.com/algebra/advanced/log/images/rules%20of%20exponents.gif <--- use rule 3 listed there, then rule 2

OpenStudy (anonymous):

\[\log \frac{ \left( x-2 \right)^3 }{\left( x-1\right)^2 }\]

OpenStudy (anonymous):

okay perfect thanks for the help guys. I am also trying to find the value of log5 (13^5). I am not sure how to calculate it or the steps involved

OpenStudy (anonymous):

Refer back to the first equation @surjithayer posted.